OpenStudy (anonymous):

Make x the subject of the formula: ax^2 + bx^2 - d^2 = p^2

OpenStudy (anonymous):

to make x as the subject u have to keep all terms containing x on oe side of the equality ax^2 + bx^2 - d^2 = p^2 ax^2 + bx^2 = p^2+d^2 or X^2(a+b)=p^2+d^2 or x^2 = (p^2+d^2)/(a+b) or x = sqrt ((p^2+d^2)/(a+b) )

OpenStudy (anonymous):

\[ax² + bx² - d² = p²\] Taking d² to the other side: \[ax²+ bx² = p²+d²\] Factorizing the L.H.S: \[ x²(a+b)=p²+d²\] Now we take (a+b) to the other side, and since its being multiplied on the L.H.S when taken to the R.H.S it will be divided: \[ x²=\frac{p²+d² }{(a+b)}\] But we only want x as the subject of the formula not x², therefore we square root the R.H.S, like so: \[ x=\sqrt{\frac{p²+d² }{(a+b)}}\]

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

No problem!(: